Problem 3- Comparing Investments Two friends deposit money into accounts over a 20-year period into an account that pays 4% compounded annually. John deposits $1000 at the end of each of 20 years. David deposits $2000 at the end of each year for 10 years, then allows the money to sit in the account to accrue more interest for the remaining 10 years. Use this information to find the following: . the amount that is in Johns account at the end of 20 years . the amount that is in Davids account after 10 years (after all the deposits are made) the amount in Davids account at the end of 20 years (10 years after the last payment is made).
Answer #1

FV = C * [(1+r)^n - 1]/i

FV = 1000*[(1+0.04)^20 - 1/0.04]

FV = 1000*(2.191123-1)/0.04

FV = 1191.123/0.04

FV in John's hand at the end of 20 years = $29778.079

AMOUNT IN DAVIDS AFTER 10 YEARS = 2000*((1+0.04)^10-1)/0.04

= 2000*0.480244/0.04

=$24012.21

Amount in David account at the end of 20 years = PV after 10 years *(1+r)^n

= 24012.21*(1+0.04)^10

= $35543.94